what is the pH of a 0.350M MgF2 solution? Ka of HF = 3.5x10^-4
Answer #1

Sol :-

Given,

Concentration of MgF2 = 0.350 M

Because, MgF2 is strong electrolyte, therefore on complete dissociation it gives 1 mol of Mg2+ ion and 2 moles of F- ions as :

MgF2 ----------------------> Mg2+ + 2F-

So,

Concentration of F- = 2 x Concentration of MgF2 = 2 x 0.350 M = 0.700 M

Also given,

Acid dissociation constant of HF = ka = 3.5 x 10-4

So,

Base dissociation constant of F- = kb = kw (Ionic product of water)/ka

kb = 1.0 x 10-14 / 3.5 x 10-4

kb = 2.86 x 10-11  

ICE table of F- is :

................................F-..................+.................H2O <-----------------------> HF................+...............OH-

Initial (I)..................0.700 M.....................................................................0.0 M.............................0.0 M

Change (C)..............-y...............................................................................+y..................................+y

Equilibrium (E)......(0.700-y) M....................................................................y M................................y M

y = Amount dissociated per mole

Expression of kb is :

kb = [HF].[OH-] / [F-]

2.86 x 10-11 = y2/((0.700-y)

y<<<0.700, so neglect y as compare to 0.700

y2 = 2.86 x 10-11 x 0.700

y = 4.47 x 10-6

So,

[OH-] = y M = 4.47 x 10-6  M

Now,

pOH = - log [OH-]

So,

pOH = - log 4.47 x 10-6 = 5.35

Also,

pH + pOH = 14

So,

pH = 14-pOH

pH = 14 - 5.35

pH = 8.65

Hence, pH = 8.65